what is 11x^2=27 i dont really get how to do that
Answer by Alan3354(69443)

what is 11x^2=27 i dont really get how to do that
--------------
what is 11x^2=27
That's an equation in x.
--------------------
I'll assume you want to solve for x, you didn't say.
11x^2=27
Divide thru by 11
x^2 = 27/11
x = +sqrt(27/11) =~ 1.5667
x = -sqrt(27/11) =~ -1.5667